$SPAD/src/input table.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool table.output
)set message test on
)set message auto off
)clear all
--S 1 of 18
t: Table(Polynomial Integer, String) := table()
--R
--R
--R   (1)  table()
--R                                      Type: Table(Polynomial(Integer),String)
--E 1
--S 2 of 18
setelt(t, x**2 - 1, "Easy to factor")
--R
--R
--R   (2)  "Easy to factor"
--R                                                                 Type: String
--E 2
--S 3 of 18
t(x**3 + 1) := "Harder to factor"
--R
--R
--R   (3)  "Harder to factor"
--R                                                                 Type: String
--E 3
--S 4 of 18
t(x)        := "The easiest to factor"
--R
--R
--R   (4)  "The easiest to factor"
--R                                                                 Type: String
--E 4
--S 5 of 18
elt(t, x)
--R

--R
--R   (5)  "The easiest to factor"
--R                                                                 Type: String
--E 5
--S 6 of 18
t.x
--R
--R
--R   (6)  "The easiest to factor"
--R                                                                 Type: String
--E 6
--S 7 of 18
t x
--R
--R
--R   (7)  "The easiest to factor"
--R                                                                 Type: String
--E 7
--S 8 of 18
t.(x**2 - 1)
--R
--R
--R   (8)  "Easy to factor"
--R                                                                 Type: String
--E 8
--S 9 of 18
t (x**3 + 1)
--R
--R
--R   (9)  "Harder to factor"
--R                                                                 Type: String
--E 9
--S 10 of 18
keys t
--R
--R
--R             3      2
--R   (10)  [x,x  + 1,x  - 1]
--R                                              Type: List(Polynomial(Integer))
--E 10
--S 11 of 18
search(x, t)
--R
--R

--R   (11)  "The easiest to factor"
--R                                                      Type: Union(String,...)
--E 11
--S 12 of 18
search(x**2, t)
--R
--R
--R   (12)  "failed"
--R                                                    Type: Union("failed",...)
--E 12
--S 13 of 18
search(x**2, t) case "failed"
--R
--R
--R   (13)  true
--R                                                                Type: Boolean
--E 13
--S 14 of 18
remove!(x**2-1, t)
--R
--R
--R   (14)  "Easy to factor"
--R                                                      Type: Union(String,...)
--E 14
--S 15 of 18
remove!(x-1, t)
--R
--R
--R   (15)  "failed"
--R                                                    Type: Union("failed",...)
--E 15
--S 16 of 18
#t
--R
--R
--R   (16)  2
--R                                                        Type: PositiveInteger
--E 16
--S 17 of 18
members t
--R
--R
--R   (17)  ["The easiest to factor","Harder to factor"]
--R                                                           Type: List(String)

--E 17
--S 18 of 18
count(s: String +-> prefix?("Hard", s), t)
--R
--R
--R   (18)  1
--R                                                        Type: PositiveInteger
--E 18
)spool
)lisp (bye)

———-

References^{[1] nothing}

7